OpenStudy (anonymous):

find area of trainagle 1 1/4 is base and height is 4/9

OpenStudy (cwrw238):

area = half of the base * the height

OpenStudy (anonymous):

i got 1 1/9 is that correct

OpenStudy (anonymous):

No, that's not right. I'm working it out right now

OpenStudy (anonymous):

It's equivalent to .277777778

OpenStudy (anonymous):

how did you get that

OpenStudy (phi):

Can you change the base to an improper fraction?

OpenStudy (anonymous):

half of 1 and 1/4 is .625. and then multiply that by 4/9 and you get .27777778

OpenStudy (anonymous):

phi yes you can

OpenStudy (phi):

The formula for the area of a triangle is \[ A= \frac{1}{2} b\cdot h\] where b is the base and h is the height (or altitude) the easiest way is to change the base \(= 1 \frac{1}{4}\) to an improper fraction then multiply the fractions b*h then multiply by 1/2

OpenStudy (phi):

I know I can. Can you? if so what is it?

OpenStudy (anonymous):

5/4

OpenStudy (phi):

now multiply the fractions \[ \frac{1}{2}\cdot\frac{5}{4}\cdot \frac{4}{9} \] multiply tops, and multiply bottoms (but you can simplify by first cancelling the 4's)

OpenStudy (anonymous):

5/4 times 4/9= 20/36 then do i multiply by half

OpenStudy (phi):

yes

OpenStudy (anonymous):

is its 5/18

OpenStudy (phi):

but notice it is easier to first cancel the 4's You can do that because order does not matter. \[ \frac{1}{2}\cdot\frac{\cancel{4}}{\cancel{4}}\cdot \frac{5}{9} = \frac{5}{18}\]

OpenStudy (anonymous):

thank you